Calculate the morality of 20.5 g sodium chloride (NaCl) compound with molar mass of 58.44g/mole placed in 55.0 mL of total solution. 


1
Expert's answer
2021-03-22T08:04:17-0400

Molarity , "M= moles\/1000"

Moles = 20.5/58.4= 0.35 moles

Molarity= 0.35/1000

=3.5× 10-4M